OpenStudy (anonymous):

**Medal for good answer Directions: Use the Pythagorean Theorem to find the distance, to the nearest tenth, between each pair of points. 11. *shows a graph**Y is on (0,1) and Z is on (4,5) 12. *shows a graph**L is on (-2,4) and M is on (3,-4) Show work please I just want to know how to solve it:)

OpenStudy (anonymous):

use the pythagorean theorem to solve it, like it says. a^2 + b^2 = c^2. The line between the two points makes the hypotenuse of a right triangle with the legs being the difference in x coords and difference in y coords.

OpenStudy (anonymous):

Im sorry...wish I could give you a medal but that's just not explained enough for me... Including the answer would be great

OpenStudy (anonymous):

"I just want to know how to solve it" For the first one, the difference in x coords is 4, and the difference in y coords is 4. using the theorem: 4^2 + 4^2 = c^2 c^2 = 32 c = square root of 32

OpenStudy (dumbcow):